# From Tokens to Thoughts: How LLMs and Humans Trade Compression for Meaning

### v3

@@ 3.2 Large Language Models Under Study
Our analysis features encoder-only models from the BERT family (e.g., BERT-Large (Devlin et al., 2019; He et al., 2020; Zhuang et al., 2021)). The majority are decoder-only autoregressive models, including: six Llama family models (1B to 70B, e.g., Llama 3.1 70B (Touvron et al., 2023a, b; Grattafiori et al., 2024)); five Gemma family models (2B to 27B (Team et al., 2024, 2025)); thirteen Qwen family models (0.5B to 72B (Bai et al., 2023; Yang et al., 2024)); four Phi family models (e.g., Phi-4 (Javaheripi et al., 2023; Abdin et al., 2024; Abouelenin et al., 2025)); and a Mistral 7B model (Karamcheti et al., 2021). Appendix A.2 provides a comprehensive list of all model variants, identifiers, and architectural details.
@@ 4.2 The $\mathcal{L}$ Objective: Balancing Representational Complexity and Semantic Distortion
The Complexity (Rate) Term: The first component, Complexity($X,C$), measures the informational cost or intricacy of representing the original items $X$ through their assignments to clusters $C$. It is quantified by the mutual information $I(X;C)$ between the items and their cluster labels. A lower $I(X;C)$ signifies greater compression, meaning the cluster assignments $C$ make the specific items $X$ more predictable (i.e., require less information to specify beyond the cluster label). Defining $I(X;C)=H(X)-H(X|C)$, and assuming $|X|$ equiprobable unique items for the initial entropy calculation ($H(X)=\log_{2}|X|$), the conditional entropy is $H(X|C)=\frac{1}{|X|}\sum_{c\in C}|C_{c}|\log_{2}|C_{c}|$. This assumes that for this complexity calculation, items within each cluster $C_{c}$ (of size $|C_{c}|$) are indistinguishable beyond their shared label $c$. Thus:
@@ 5 Unpacking Representational Strategies: An Empirical Investigation
Experimental Recap: LLM token embeddings from our benchmark datasets (Rosch, 1973a, 1975; McCloskey and Glucksberg, 1978) were clustered (k-means; $K$ matching human category counts). Alignment with human categories was measured using AMI, NMI, and ARI (AMI shown in Figure 1; see Appendices A.3, A.4 for full details).
Results and Observations: Across all tested LLMs, derived conceptual clusters aligned with human categories significantly above random chance (Figure 1, showing averaged AMI scores). This indicates their semantic spaces encode information supporting human-like grouping at a macro level. Notably, the BERT family (especially BERT-large-uncased) demonstrated robust alignment, often comparable to or exceeding that of much larger decoder-only models. This suggests that architectural or pre-training factors, not just scale, influence the formation of human-like categorical structures.
Figure 1: LLM-derived Clusters Show Above-Chance Alignment with Human Conceptual Categories. Adjusted Mutual Information (AMI) between human categories and LLM-embedding clusters versus model size. Results are averaged over three psychological datasets. All models perform significantly better than random clustering. BERT’s performance is notably strong.
@@
Experimental Recap: For this RQ, as detailed in this section’s introduction, we compared human typicality judgments from the cognitive science datasets (Rosch, 1973a, 1975; McCloskey and Glucksberg, 1978) with an LLM-based measure. Specifically, we calculated the cosine similarity between each item’s token embedding and the token embedding of its *human-assigned category name* (e.g., ‘robin‘ vs. ‘bird‘). These item-to-category-label similarities were then correlated (Spearman’s $\rho$ (Wissler, 1905)) with human-rated typicality scores.
Results and Observations: Spearman correlations between LLM-derived item-to-category-label similarities and human typicality judgments are generally modest across most models and datasets (Table 2 in Appendix A.5; Figure 6). Although some correlations reach statistical significance ($p<0.05$), their magnitudes typically indicate a limited correspondence. This pattern suggests that items humans perceive as highly typical of a category are not consistently represented by LLMs as substantially more similar to that category label’s embedding. While BERT-large-uncased occasionally exhibited slightly stronger correlations, these remained moderate (Table 2). Consequently, no tested model robustly replicated the full spectrum of human typicality gradients using this measure. Appendix A.6 provides further visualizations supporting these observations.
@@
Interpretation: These findings suggest that while LLMs can identify features for broad categorization, their organization of semantic space around explicit category labels does not fully mirror the nuanced prototype structures evident in human typicality judgments. The factors driving an item’s embedding similarity to its category label’s embedding in LLMs may differ from the rich, multifaceted criteria (e.g., perceptual attributes, functional roles) underpinning human typicality. LLMs might instead capture a more statistically uniform association to category labels, thereby under-representing the graded, prototype-centric nature of human concepts. This divergence in capturing fine-grained semantics leads to our subsequent inquiry into overall information processing efficiency.
@@
(a) Human conceptual categories exhibit higher mean entropy. Mean cluster entropy ($S_{\alpha}$) vs. number of clusters ($K$) for LLMs and human categories (fixed $K$). Higher entropy indicates less compression.
(b) LLMs achieve more optimal $\mathcal{L}$ trade-off. Our information-theoretic objective ($\mathcal{L}$) vs. $K$. Lower $\mathcal{L}$ indicates a more statistically optimal compression-meaning balance.
Figure 2: LLMs Show More Statistically “Optimal” Compression Than Humans in Cluster Entropy and the $\mathcal{L}$ Measure. (a) Mean cluster entropy as a function of $K$ used for k-means clustering. (b) IB-RDT objective ($\mathcal{L}$) as a function of $K$ used for k-means clustering. Human categories consistently show higher entropy and $\mathcal{L}$ values. Results shown for Rosch (1975) dataset; full results in Appendix A.8.
@@ Appendix A Limitations
  • We focus on static, context-free representations. LLMs may fall short in capturing context sensitivity, as human concepts are influenced by factors beyond raw compression efficiency (experience, social interaction, cultural context).
@@
Future work could address these by expanding to other languages, exploring alternative cognitive models, dynamic representations, and testing these principles on different architectures or in real-world applications.
@@ A.2 LLM Details
  • QWEN family: qwen2-0.5b, qwen2.5-0.5b, qwen1.5-0.5b, qwen2.5-1.5b, qwen5-1.5b, qwen1.5-1.5b, qwen1.5-4b, qwen2.5-4b, qwen2-7b, qwen1.5-14b, qwen1.5-32b, qwen1.5-72b [Bai et al., 2023, Yang et al., 2024].
@@
  • Gemma family: gemma-2b, gemma-2-2b, gemma-7b, gemma-2-9b, gemma-2-27b [Team et al., 2024, 2025].
@@
A.3 Additional Clustering Metrics

### v4

@@ 3.1 Human Conceptual Baselines: Empirical Data from Seminal Cognitive Science
Appendix  A.6 complements this benchmark by reporting WordNet-based polysemy statistics for every item. These figures characterize the lexical ambiguity of the stimuli and frame the subsequent embedding analyses. The analysis shows that polysemy is very rare in our data and thus cannot explain the findings we report later in the paper.
@@ 3.2 Large Language Models Under Study
Our analysis features encoder-only models from the BERT family (e.g., BERT-Large (Devlin et al., 2019; He et al., 2020; Zhuang et al., 2021)). The majority are decoder-only autoregressive models, including: six Llama family models (1B to 70B, e.g., Llama 3.1 70B (Touvron et al., 2023a, b; Grattafiori et al., 2024)); five Gemma family models (2B to 27B (Team et al., 2024, 2025)); thirteen Qwen family models (0.5B to 72B (Bai et al., 2023; Yang et al., 2024)); four Phi family models (e.g., Phi-4 (Javaheripi et al., 2023; Abdin et al., 2024; Abouelenin et al., 2025)); a Mistral 7B model (Karamcheti et al., 2021); two GPT-2 models of different scales (Radford et al., 2019); and checkpoints from the OLMo-7B model (Groeneveld et al., 2024). Other, less common architectures include a single encoder-decoder T5 model (Raffel et al., 2020), two text encoders from CLIP ViT (ViT-B/32 and ViT-B/16) (Radford et al., 2021), and the classic encoder models GloVe (Pennington et al., 2014) and Word2Vec (Mikolov et al., 2013a, b). Note: GloVe and Word2Vec provide static embeddings only, and thus their analysis is constrained. Appendix A.2 provides a comprehensive list of all model variants, identifiers, and architectural details.
@@
To complement the analysis, we extend our methodology beyond static embeddings to contextual representations. For each LLM, we extract token-level embeddings using a set of carefully controlled prompts (see Appendix A.3), enabling a principled basis for directly comparing static token representations to their contextualized counterparts. This allows for a similar assessment while mitigating confounding issues such as prompt length and token position. We found minimal variation in results across prompt templates and pooling methods. Appendix A.3 provides a detailed exploration.
Finally, tokenizer-overlap analysis (Appendix A.7) shows substantial lexical sharing across model families; neither token count nor tokenization overlap correlates with AMI, confirming that tokenizer effects are negligible and thus cannot explain our findings.
@@ 4.2 The $\mathcal{L}$ Objective: Balancing Representational Complexity and Semantic Distortion
The Complexity (Rate) Term: The first component, Complexity($X,C$), measures the informational cost or intricacy of representing the original items $X$ through their assignments to clusters $C$. It is quantified by the mutual information $I(X;C)$ between the items and their cluster labels. A lower $I(X;C)$ signifies greater compression, meaning the cluster assignments $C$ make the specific items $X$ more predictable (i.e., require less information to specify beyond the cluster label). Defining I(X;C)=H(X)H(X|C)I(X;C)=H(X)-H(X|C), and assuming $|X|$ equiprobable unique items for the initial entropy calculation ($H(X)=\log_{2}|X|$), the conditional entropy is $H(X|C)=\frac{1}{|X|}\sum_{c\in C}|C_{c}|\log_{2}|C_{c}|$. This assumes that for this complexity calculation, items within each cluster $C_{c}$ (of size $|C_{c}|$) are indistinguishable beyond their shared label $c$. Thus:
@@ 5 Unpacking Representational Strategies: An Empirical Investigation
Experimental Recap: LLM token embeddings from our benchmark datasets (Rosch, 1973a, 1975; McCloskey and Glucksberg, 1978) were clustered (k-means; $K$ matching human category counts). Alignment with human categories was measured using AMI, NMI, and ARI (AMI shown in Figure 1; see Appendices A.8, A.9 for full details). Both the static and the contextual embeddings were extracted. The contextual embeddings were measured layer-wise, and reported using Peak AMI across the layers. Note that GloVe and Word2Vec only have static embeddings.
Results and Observations: Across all tested LLMs, derived conceptual clusters aligned with human categories significantly above random chance (Figure 1, showing averaged peak AMI scores). This indicates their semantic spaces encode information supporting human-like grouping at a macro level. Notably, the BERT family (especially BERT-large-uncased) demonstrated robust alignment, often comparable to or exceeding that of much larger decoder-only models. This suggests that architectural or pre-training factors, not just scale, influence the formation of human-like categorical structures.
Figure 1: LLM-derived Clusters Show Above-Chance Alignment with Human Conceptual Categories. Adjusted Mutual Information (AMI) between human categories and LLM-embedding clusters versus model size. Results are averaged over three psychological datasets. All models perform significantly better than random clustering. Encoder architectures achieve consistently high performance across modalities and scales; notably, the classic Word2Vec and GloVe models, despite offering only static, outdated embeddings, reach static AMI scores that match the peak AMI of modern LLMs. See Appendix A.4 for a comparison of static, mean, and peak AMI.
@@
Influence of Training Dynamics on Alignment: To explore how semantic structure develops during training, we applied both the static embedding and the contextual (layer-wise) analysis on 10% of OLMo’s checkpoints (spaced evenly across training steps). We find that concept representations improve steadily during training, with a rapid gain in structure early on, followed by a slower, more gradual refinement. Interestingly, the model exhibits a two-phase dynamic: initial optimization moves semantic processing from deep layers to mid-network layers during rapid learning, followed by continued architectural reorganization that optimizes efficiency while maintaining performance (Appendix A.5 Figure 8). This suggests that concepts emerge early in training but are progressively consolidated and reallocated across layers, reflecting a shift from brute-force memorization toward more efficient, distributed representations. Such dynamics highlight that alignment is not just a final-state property of the trained model, but is deeply shaped by the trajectory of learning itself.
@@
Experimental Recap: For this RQ, as detailed in this section’s introduction, we compared human typicality judgments from the cognitive science datasets (Rosch, 1973a, 1975; McCloskey and Glucksberg, 1978) with an LLM-based measure. Specifically, we calculated the cosine similarity between each item’s token embedding and the token embedding of its human-assigned category name (e.g., ‘robin‘ vs. ‘bird‘). These item-to-category-label similarities were then correlated (Spearman’s $\rho$ (Wissler, 1905)) with human-rated typicality scores.
We employed two analysis approaches: static-layer analysis utilized static items/category embeddings from the embedding layer, while peak AMI layer analysis utilized items/category contextual embeddings from the layer that maximized AMI as measured in RQ1. To extract category embeddings in the peak AMI approach, we replaced the item with the category name in the same prompt template (see template and template- and pooling robustness analysis in Appendix A.4).
Results and Observations: Spearman correlations between LLM-derived item-to-category-label similarities and human typicality judgments are generally modest across most models and datasets, with peak AMI layers showing particularly weak correlations (Tables 3,  4 in Appendix A.10; Figure 15). Even when correlations reach statistical significance ($p<0.05$), their magnitudes typically indicate a limited correspondence.
Several architectural patterns emerge: representation-focused models (Word2Vec, GloVe) and some encoder models (T5-large, the two ViT encoders, and BERT large) demonstrate stronger static-layer performance than decoder-only models (Llama, Gemma, etc.), suggesting that models trained specifically for representation learning may be more effective for capturing semantic category relationships in their embeddings. In contrast, modern decoder-only models show consistently lower static-layer correlations, potentially due to their training objectives being more focused on next-token prediction rather than semantic representation.
Notably, peak AMI layer analysis shows that correlations weaken substantially relative to static-layer analysis across most architectural variants, suggesting that cross-cluster relations (RQ1) and within-cluster organization (RQ2) are encoded at different layers. This implies that downstream applications should select layers based on the resolution they require; whether capturing broad cross-cluster structure or fine-grained within-cluster organization. Furthermore, while some architectural approaches capture certain aspects of human typicality judgments, their internal representations exhibit only modest alignment with human-perceived fine-grained semantic distinctions, such as item typicality or psychological distance to category prototypes. As a result, no single architectural approach robustly reproduces the full spectrum of human typicality gradients using this measure. Additional visualizations supporting these findings are provided in Appendix A.11.
@@
Interpretation: These findings suggest that while LLMs can identify features for broad categorization, their organization of semantic space around explicit category labels does not fully mirror the nuanced prototype structures evident in human typicality judgments. The factors driving an item’s embedding similarity to its category label’s embedding in LLMs may differ from the rich, multifaceted criteria (e.g., perceptual attributes, functional roles) underpinning human typicality. LLMs might instead capture a more statistically uniform association to category labels, thereby under-representing the graded, prototype-centric nature of human concepts. The tension between representational compactness and semantic fidelity becomes particularly evident when comparing peak AMI layers to static layers: while peak AMI layers optimize the Information Bottleneck trade-off for efficient clustering, they may prioritize compression over preserving the nuanced semantic information needed for human-like typicality judgments. This divergence in capturing fine-grained semantics leads to our subsequent inquiry into overall information processing efficiency.
@@
(a) Distortion as a function of Complexity demonstrates the balance of the Information Bottleneck in our information-theoretic objective ($\mathcal{L}$). The plot illustrates how the IB balances compression (complexity) against information preservation (distortion). Encoder models (BERT, ViT, classic models) show lower distortion rates compared to all the decoder models across all complexity values.
(b) Contextual embeddings achieve better-than-human $\mathcal{L}$ trade-off. Our information-theoretic objective ($\mathcal{L}$) vs. $K$. Lower $\mathcal{L}$ indicates a more statistically optimal compression-meaning balance compared to human baseline.
Figure 2: LLMs Show More Statistically “Optimal” Compression Than Humans in the $\mathcal{L}$ Measure. (a) Distortion as a function of Complexity showing the trade-off relationship between the two terms in the IB-RDT objective ($\mathcal{L}$). Encoder models (BERT, ViT, classic models) show lower distortion rates compared to all the decoder models across all complexity values. (b) IB-RDT objective ($\mathcal{L}$) as a function of $K$ used for k-means clustering using contextual embeddings. Lower $\mathcal{L}$ indicates a more statistically optimal compression-meaning balance. Human categories consistently show higher $\mathcal{L}$ values. Results shown for Rosch (1975) dataset; full results in Appendix A.13.
@@
The Objective $\mathcal{L}$ vs. Downstream Task Performance: The relationship between concept compression, as measured by the $\mathcal{L}$ objective function, and downstream task performance shows no correlation, indicating that concept compression captures an orthogonal property to benchmark accuracy. See Appendix A.14 for detailed analysis.
@@ Appendix A Limitations
Future work could address these by expanding to other languages, exploring alternative cognitive models, and testing these principles on different architectures or in real-world applications.
@@ A.2 LLM Details
  • QWEN family: qwen2-0.5b, qwen2.5-0.5b, qwen1.5-0.5b, qwen2.5-1.5b, qwen2-1.5b, qwen1.5-1.5b, qwen1.5-4b, qwen2.5-4b, qwen2-7b, qwen1.5-14b, qwen1.5-32b, qwen1.5-72b, qwen2.5-72b [Bai et al., 2023, Yang et al., 2024].
@@
  • Gemma family: gemma-2b, gemma-2-2b, gemma-7b, gemma-2-9b [Team et al., 2024, 2025].
@@
  • GPT family: gpt2, gpt2-medium [Radford et al., 2019].
  • OLMo family: olmo-7b [Groeneveld et al., 2024].
  • ViT family: Clip ViT-B/32, Clip ViT-B/16 [Radford et al., 2021].
  • Classic static embeddings: GloVe, Word2Vec [Pennington et al., 2014, Mikolov et al., 2013a, b].
A.3 Contextual Prompts and Pooling Strategies
Contextual embeddings of LLMs require feeding words into the model through a prompt. Because tokenizers often split a word into multiple tokens, and since some items in our datasets consist of two or more words, we face a design choice regarding how to aggregate token representations. In our methodology, we adopt average pooling over the actual tokens, ensuring that all subword pieces contribute equally. Figures 3 and 4 reveal that the average pooling strategy achieves consistent performance and demonstrates the tightest distribution, making it the most reliable choice for our research.
For prompts, we selected a neutral template, "This is a {word}. " (with a trailing space), designed to minimize any additional semantic bias on the target item. Figures 5 and 6 show this prompt to balance performance and consistency, making it ideal for baseline comparisons.
In this section, we explore alternative pooling strategies and evaluate a diverse set of prompt templates across multiple models.
Figure 3: Average pooling demonstrates consistent performance across different models, making it the most reliable choice for future research. Each point corresponds to a prompt template applied to a model.
Pooling Strategies.
We compare four common approaches:
  • Avg - mean over all tokens representing the word
  • First - representation of the first token
  • Last - representation of the last token
  • Sentence Avg - mean over the full sentence embedding
Prompt Templates.
To test robustness, we design eight templates spanning different linguistic framings:
  • "This is a {word}."
  • "This is a {word}. " (with trailing space)
  • "The concept of {word} is"
  • "When we think of {word}, we consider"
  • "A typical {word} would be"
  • "Examples of {word} include"
  • "The category {word} contains"
  • "One kind of {word} is"
Models.
We evaluate eight representative LLMs covering major architectures:
  • bert-large-uncased (BERT family)
  • deberta-large (DeBERTa family)
  • gemma-2-2b (Gemma family)
  • Llama-3.2-1B (Llama family)
  • Mistral-7B-v0.3 (Mistral family)
  • phi-2 (Phi family)
  • Qwen2.5-1.5B (Qwen family)
  • roberta-large (RoBERTa family)
Figure 4: Average pooling demonstrates the tightest distribution, indicating the highest consistency and reliability across different conditions. Performance Distribution by Pooling Strategy - Box plots showing AMI distribution for each pooling strategy
Figure 5: The neutral prompt template "This is a word. " demonstrates balanced performance and moderate consistency, making it ideal for baseline comparisons. Performance Distribution across various prompts.
Figure 6: The neutral template "This is a {word}. " remains a stable choice across families. Heatmap showing template performance per model.
A.4 Static vs. Contextual AMI Exploration
To understand how LLMs develop conceptual alignment with human categories, we examine the progression from static to contextual embeddings. Figure 7 presents three complementary views of this progression across different model scales and architectures.
The left subplot shows Static AMI scores, which represent the conceptual alignment achieved by models’ input embeddings before any contextual processing (i.e., the E matrix embeddings of the target word). These scores reveal that even at the most basic level, LLMs encode semantic information that supports human-like categorical grouping. Remarkably, classic static models like Word2Vec and GloVe achieve static AMI scores that rival the peak contextual performance of modern LLMs, suggesting that fundamental conceptual structure is captured early in the learning process.
The middle subplot displays Average AMI across all layers, providing a measure of overall semantic representation quality throughout the network. This metric shows the typical performance a model achieves across its entire depth, offering insight into how consistently different layers maintain conceptual alignment. The improvement from static to average AMI demonstrates that contextual processing generally enhances rather than diminishes semantic understanding.
The right subplot reveals Peak AMI, representing the optimal conceptual alignment achieved by any single layer. This metric identifies where in the network conceptual understanding is maximized, typically occurring in middle-to-late layers before declining in the final layers. The progression from static to peak AMI shows that contextual processing not only preserves but significantly enhances the conceptual alignment present in static embeddings.
Several key insights emerge from this multi-metric analysis. First, all models demonstrate above-chance alignment even in their static embeddings, confirming that basic semantic structure is a fundamental property of learned representations. Second, the consistent improvement from static to peak AMI across all model types suggests that contextual processing universally enhances conceptual understanding rather than creating it de novo. Third, encoder architectures of different types (BERT, ViT encoders, and classic models) achieve comparable or superior performance to much larger decoder models, highlighting that architectural factors and pre-training objectives significantly influence conceptual alignment quality beyond mere model scale.
This analysis complements the main text findings by showing that LLMs do not simply achieve above-chance alignment with human categories but rather so through a systematic progression from basic to sophisticated conceptual representations, with contextual processing serving as an amplifier rather than a generator of semantic understanding.
Figure 7: LLMs begin with basic conceptual alignment in static embeddings and achieve progressively stronger alignment through contextual processing. Three subplots showing model size (log scale) versus different AMI metrics. Left: Static AMI reveals baseline categorical structure. Middle: Average AMI across layers shows overall semantic quality. Right: Peak AMI demonstrates high conceptual alignment. The consistent improvement from static to peak AMI across all model types reveals that contextual processing enhances rather than creates conceptual understanding, with encoders (BERT, ViT encoders and classic models) achieving comparable or superior performance to much larger decoder models.
A.5 Training Dynamics
The OLMo analysis examines how semantic structure develops during training by analyzing 57 intermediate checkpoints from the OLMo-7B model, representing evenly spaced sampling (every 10K training steps) spanning from 1K to 557K steps (covering approximately 4B to 2.5T tokens).
The analysis employs two complementary sampling strategies: representative sampling (6 checkpoints) captures major developmental phases at 1K, 101K, 201K, 301K, 401K, and 501K steps, while high-resolution sampling (57 checkpoints) reveals the inherent noise and fluctuations in training. Despite significant training noise, the overall semantic development follows a stable, predictable pattern captured by the representative sampling, as shown in Figure 8. The complete training trajectory with all 57 checkpoints is presented in Figure 9.
Figure 8: Left: OLMo-7B representations steadily strengthen during training: Concept representations develop rapidly at early steps, then refine more gradually over time. Right: Semantic processing shifts from deep to mid-network layers: The model undergoes a two-phase dynamic - initially moving semantic processing upward during rapid learning, then reorganizing to optimize efficiency while preserving performance. To improve readability, we present six representative checkpoints that capture the trend.
Figure 9: Complete OLMo-7B training trajectory across 57 checkpoints: This high-resolution view reveals the inherent noise and fluctuations in training, with individual checkpoint measurements varying throughout the process. Despite this variability, the overall trend aligns with the stable pattern shown in Figure 8, demonstrating that representative sampling effectively captures the underlying semantic development trajectory while filtering out training noise.
A.6 Datasets Polysemy
Scope. We quantify lexical ambiguity in our psycholinguistic stimuli by counting the distinct WordNet synsets associated with each lemma. This polysemy score lets us estimate how many alternative senses a model must implicitly conflate when it produces a single embedding for a word.
Why it matters. Consider bat, which can denote either a flying mammal or a piece of sports equipment. The same vector must account for both senses. Aggregating semantically distant senses can blur the representation and thus confound model-human comparisons, especially in tasks that rely on fine-grained semantic similarity. Explicitly tracking polysemy allows us to verify that any performance effects we observe are not artefacts of lexical ambiguity.
Results Figure 10 shows the distribution of polysemy scores. The majority of items are unambiguous (1-2 senses), but a heavy-tailed minority (e.g. Running (52 senses), Saw (28), Block (28)) is highly polysemous. This suggests that our findings are due to real differences between the models rather than polysemy-related artifacts. An additional 141 lemmas that are not in WordNet were omitted.
][t] Refer to caption
Figure 10: Polysemy is not likely to influence our results as most of the words in our dataset are concrete nouns, which tend to be less polysemous. Histogram of WordNet sense counts for the 943 lemmas in our benchmark. The dashed lines indicate the median and mean (2 senses, 3.47 senses, respectively).
A.7 Tokenizer Analysis
Rational. The tokenizer of a model has a significant influence over the representations: segmentation rules (WordPiece vs. BPE), vocabulary size and special control tokens can inflate sequence length, skew frequency statistics, and shape error patterns. To ensure fair cross-model comparisons, we therefore (i) cluster checkpoints by the tokenizer they use and (ii) quantify how much those tokenizers overlap when applied to our datasets.
Procedure. Before computing overlap, we normalize the vocabulary, stripping tokenizer-specific characters; SentencePiece prefixes (_), GPT-style BPE space prefixes (Ġ/ġ) and newline markers (Ċ), WordPiece continuations markers (##), and related block characters. After this cleanup, tokens differing only by such prefixes collapse to a shared canonical form (e.g. _house, Ġhouse, and ##house all become house). We then compute pair-wise Jaccard similarity on these cleaned vocabularies.
Table 1 summarizes the core statistics and information regarding the tokenizer types, while Figure 11 visualizes the resulting pairwise vocabulary overlap.
Findings. We find that most tokenizer families share substantial lexical overlap, often exceeding  $60\%$, suggesting a de-facto common token inventory across recent open-source models. First-generation BERT WordPiece (bert-large-uncased and bert-base-uncased) are an outlier, sharing under $16\%$ of tokens with any other group.
Model Family Mean Tokens / Item Vocabulary Size Tokenizer Type
BERT 1.65 30K WordPiece
DeBERTa & RoBERTa 2.30 50K WordPiece
GPT 2.30 50K BPE
Gemma 1.65 256K SentencePiece (subword)
Llama 2.19 128K BPE (SentencePiece/Tiktoken)
Mistral 2.35 32K BPE + control tokens
Phi 2.30 32K BPE (SentencePiece/Tiktoken)
Qwen 2.19 151.6K BPE
Table 1: Tokenizer statistics by tokenizer family. Mean Tokens/Items refer to the average of the tokens per item in our datasets. The columns Vocabulary Size and Tokenizer Type are properties of the tokenizer.
Figure 11: Substantial lexical overlap suggests that tokenization differences alone cannot explain the observed performance variations in our experiments. Vocabulary overlap between different tokenizers; most tokenizer families share substantial lexical overlap.
Model Clustering By Tokenizer Family
  • Llama: Llama-3.2-1B (representative), Llama-3.1-8B, Meta-Llama-3-8B, Llama-3.1-70B, Meta-Llama-3-70B
  • Gemma: gemma-2-9b (representative), gemma-7b, gemma-2b, gemma-2-2b
  • Mistral: Mistral-7B-v0.3 (unique)
  • Phi: phi-2 (representative), phi-1, phi-1.5
  • RoBERTa: roberta-large (unique)
  • DeBERTa: deberta-large (unique)
  • GPT: gpt2-medium (representative), gpt2
  • BERT: bert-large-uncased (representative), bert-base-uncased
  • Qwen: Qwen1.5-0.5B (representative), Qwen1.5-1.8B, Qwen1.5-14B, Qwen2-0.5B, Qwen2-7B, Qwen2.5-0.5B, Qwen2.5-1.5B, Qwen2.5-32B, Qwen1.5-32B, phi-4111phi-4 has a different tokenizer than the rest of Phi family. The results of its tokenizer match the tokenizer of the Qwen family.
A.8 Additional Clustering Metrics